# Intrinsic Heisenberg-type lower bounds on spacelike hypersurfaces in general relativity

### v2

@@ 2.2 Momentum operator and variance decomposition
In an $h$-orthonormal frame the (DeWitt–symmetrized) momentum components are
$P_{a}:=-\,i\hbar\Big(\nabla^{h}_{X_{a}}+\tfrac{1}{2}\,\mathrm{div}_{h}X_{a}\Big),$ (2)
which are symmetric on the natural Dirichlet form domain (e.g. $H^{1}_{0}(\Omega)$ on a domain $\Omega\subset\Sigma$). The momentum variance with respect to $\psi$ is the quadratic form
$\mathrm{Var}_{\psi}(P):=\Big\langle\big\|P-\langle P\rangle\big\|_{h}^{2}\Big\rangle$ (3)
with $P:=(P_{a})_{a=1}^{n}$ and $\langle P\rangle=(\langle P_{a}\rangle)_{a=1}^{n}$. For clarity, we set $\big\langle\|P\|_{h}^{2}\big\rangle:=\sum_{a=1}^{n}\langle\psi,\,P_{a}^{\dagger}P_{a}\,\psi\rangle$; hence (by Green’s formula with Dirichlet data), we obtain the index‑free identity
$\big\langle\,\|P\|^{2}_{h}\,\big\rangle=-\,\hbar^{2}\langle\psi,\,\Delta_{h}\psi\rangle=\hbar^{2}\,\|\nabla^{h}\psi\|_{L^{2}}^{2},$
which leads to
$\mathrm{Var}_{\psi}(P)=\hbar^{2}\,\|\nabla^{h}\psi\|_{L^{2}}^{2}\;-\;\|\langle P\rangle\|^{2}_{h}.$ (4)
@@
We measure momentum uncertainty by the standard deviation
@@
Geometric content. Eqs. (2)–(5) are framed entirely in the induced Riemannian geometry $(\Sigma,h)$: they are coordinate- and foliation-independent, use only bulk Dirichlet data on $B_{\Sigma}(p,r)$, and prepare the ground for comparisons with the Dirichlet spectrum via Rayleigh–Ritz.
From a physicist’s viewpoint, (2)–(5) say: “variance of momentum on a curved, bounded region equals kinetic energy (in units $\hbar^{2}$) with the global drift subtracted.” Because (4) depends only on the intrinsic geometry of $(\Sigma,h)$ via $\nabla^{h}$ and $\mu_{h}$, it is coordinate-independent and well adapted to comparisons with spectral data.
@@ 3 An intrinsic Heisenberg-type lower bound
The identity (4) – i.e. the variance decomposition $\mathrm{Var}_{\psi}(P)=\hbar^{2}\,\|\nabla^{h}\psi\|_{L^{2}}^{2}-\|\langle P\rangle\|_{h}^{2}$ – is derived in the Dirichlet setting with $\psi\in H^{1}_{0}(B_{\Sigma}(p,r))$, where Green’s formula yields $\langle\psi,P^{\dagger}P\,\psi\rangle=\hbar^{2}\|\nabla^{h}\psi\|_{L^{2}}^{2}$. For Neumann data ($\partial_{n}\psi=0$) the boundary term in Green’s identity also vanishes, so the same representation holds on the Neumann form domain $H^{1}(B_{\Sigma}(p,r))$; however, the first Neumann eigenvalue is 0 on connected domains, which makes our spectral lower bound degenerate. For Robin data ($\partial_{n}\psi+\kappa\psi=0$) a surface contribution proportional to $\int_{\partial B_{\Sigma}(p,r)}|\psi|^{2}$ survives, so the variance cannot be expressed purely by the bulk Dirichlet form. Accordingly, in the following we impose Dirichlet boundary conditions throughout and use the Dirichlet Laplacian with its Rayleigh–Ritz characterization. Write $\lambda_{1}(\Omega;h)$ for the first Dirichlet eigenvalue of $\Delta_{h}$ on a bounded domain $\Omega\subset\Sigma$:
@@
$\sigma_{p}(\psi)\;\geq\;\hbar\,\sqrt{\lambda_{1}\!\left(B_{\Sigma}(p,r);h\right)}.$ (7)
Equality holds if and only if $|\psi|$ is a first Dirichlet eigenfunction of $\Delta_{h}$ on $B_{\Sigma}(p,r)$ and $\nabla^{h}(\arg\psi)$ is ($\nu$-a.e.) constant on $B_{\Sigma}(p,r)$.
@@
Using the Madelung identity $\|\nabla^{h}\psi\|_{h}^{2}=\|\nabla^{h}u\|_{h}^{2}+u^{2}\|\nabla^{h}\phi\|_{h}^{2}$, the variance formula (4) gives the exact decomposition
@@
Since $\|\psi\|_{L^{2}(B_{\Sigma}(p,r),d\mu_{h})}=1$, the weight $u^{2}\,d\mu_{h}$ is a probability measure on $B_{\Sigma}(p,r)$. Applying the Cauchy–Schwarz inequality with respect to this weight, we obtain
$\sum_{a}\!\Big(\int u^{2}X_{a}\phi\,d\mu_{h}\Big)^{\!2}\;\leq\;\int u^{2}\sum_{a}(X_{a}\phi)^{2}\,d\mu_{h}\;\equiv\;\int u^{2}\|\nabla^{h}\phi\|_{h}^{2}\,d\mu_{h}.$ (10)
@@
Invoking (1), the difference of expectations simplifies to
@@
In (14) the fluctuation term is written as a sum of variances of the scalar fields $X_{a}\phi$ with respect to the probability measure $d\nu$. One may rewrite this sum as the squared $h$-norm of a centered gradient vector and then take the $\nu$–expectation. Concretely,
@@
for $\psi=u\,e^{i\phi}$. Hence, equality in (11) holds precisely when the gradient $\nabla^{h}\phi$ is $\nu$-a.e. constant. This is the phase part of the equality condition in Theorem 3.1; together with $|\psi|$ being a first Dirichlet eigenfunction it yields equality in (7).
By the Dirichlet Rayleigh–Ritz characterization (6) (see [1, Ch. 3]), for any $0\neq u\in H^{1}_{0}(B_{\Sigma}(p,r))$,
@@
Taking $u=|\psi|$ with $\|\psi\|_{L^{2}}=1$ (hence $\|u\|_{L^{2}}=1$) and combining this with the lower bound (11), we obtain
@@
Taking square roots yields (7). $\square$
Remark 3.2 (Coordinate invariance and kinetic-energy window).
The estimate is purely intrinsic to the Riemannian data $(B_{\Sigma}(p,r),h)$: neither coordinates nor the ambient foliation play any role in its formulation or proof. Combining (4) with the exact decomposition (15) yields the kinetic-energy window
@@
Equality at the upper endpoint holds precisely when the mean momentum vanishes, $\langle P\rangle=0$; equality at the lower endpoint holds precisely when $\nabla^{h}\phi$ is $\nu$-a.e. constant. If the lower-endpoint equality holds but $|\psi|$ is not a ground-state Dirichlet eigenfunction, then the Rayleigh quotient is strict and the bound (7) is correspondingly strict: $\sigma_{p}(\psi)>\hbar\sqrt{\lambda_{1}}$.
Remark 3.3 (On the ‘constant phase gradient’ condition and frame independence).
By the exact variance decomposition (15), equality at the lower endpoint of the kinetic-energy window (18) (equivalently of (11)) holds if and only if the $\nu$-variance of the phase gradient vanishes:
@@
Realizability / integrability. Writing $V:=\sum_{a=1}^{n}c_{a}\,X_{a}$, the condition above requires $\nabla^{h}\phi=V$. Such a phase exists if and only if $d(V^{\flat})=0$ (equivalently, $\nabla_{i}V_{j}-\nabla_{j}V_{i}=0$). On curved balls this typically fails; hence equality in (7) is often not attained (the bound is sharp but strict). In the Euclidean ball $B_{r}(0)\subset\mathbb{R}^{n}$, however, $d(V^{\flat})=0$ for constant $V$, and affine phases $\phi(x)=k\!\cdot\!x$ realize the lower endpoint together with the Dirichlet ground-state modulus.
Remark 3.4 (Model saturating the bound with nonzero mean momentum).
Let $B_{r}(0)\subset\mathbb{R}^{n}$ carry the Euclidean metric $h$. Let $u_{1}>0$ denote the $L^{2}$-normalized first Dirichlet eigenfunction of $\Delta_{h}$ on $B_{r}(0)$, with eigenvalue $\lambda_{1}(B_{r}(0);h)=:\lambda_{1}$. For fixed $k\in\mathbb{R}^{n}$ set

### v4

@@ 2.2 Momentum operator and variance decomposition
Before strictly choosing the momentum operators on the curved manifold, we briefly motivate the choice of the reference frame. While the traditional DeWitt quantization [26, 27] can be consistently formulated in a natural coordinate basis $\{\partial_{i}\}$, the resulting expression for the momentum variance inevitably involves the inverse metric $h^{ij}(x)$ to contract the operator indices. This explicit metric dependence obscures the physical interpretation, as it mixes kinematic fluctuations with geometric terms in a non-trivial way. In contrast, adopting an $h$-orthonormal frame $\{X_{a}\}$ effectively diagonalizes the local geometry, allowing the variance to be expressed as a simple sum of squares analogous to the Euclidean case. This generalized approach is advantageous as it facilitates a clean, coordinate-independent separation of the pure quantum fluctuations from the intrinsic curvature effects encoded in the geometric potential derived below.
From the structural point of view, the explicit form of the momentum operators associated with a given orthonormal frame is strongly constrained by general axiomatic schemes for quantization on configuration manifolds. In Segal’s representation-theoretic approach to quantization of nonlinear systems [28] and in the framework of geometric quantization [29], one requires that configuration observables $f\in C^{\infty}_{0}(\Sigma)$ act as multiplication operators and that each momentum observable $P_{X}$ associated with a real vector field $X$ be represented by a first-order differential operator whose principal symbol coincides with the classical momentum. In addition, the commutator with configuration observables is required to reproduce the classical Poisson brackets,
$f,P_{X}]=i\hbar\,X(f),\qquad f\in C^{\infty}_{0}(B_{\Sigma}(p,r)),$ (2)
and the resulting operators must be symmetric on $L^{2}(B_{\Sigma}(p,r),d\mu_{h})$ with respect to the Riemannian volume $d\mu_{h}$. The Borel quantization programme developed in [30, 31] shows that, under these assumptions and in the presence of a fixed Riemannian structure $(\Sigma,h)$, there is a distinguished choice for $P_{X}$ that is covariant under changes of orthonormal frame and compatible with the unitary implementation of the configuration space symmetries.
Specializing to an $h$-orthonormal frame $\{X_{a}\}_{a=1}^{n}$, the corresponding momentum components are then realized as the differential operators
$P_{a}:=-i\hbar\left(\nabla^{h}_{X_{a}}+\frac{1}{2}\,\mathrm{div}_{h}X_{a}\right),\qquad a=1,\dots,n,$ (3)
which coincide with DeWitt’s covariant momenta in an orthonormal frame [26, 27] and reduce to the familiar $p_{i}=-i\hbar\,\partial_{i}$ in Euclidean space with a global Cartesian frame. In the language of [30, 31], the $\frac{1}{2}\,\mathrm{div}_{h}X_{a}$ term is precisely the correction dictated by the requirement that the map $X\mapsto P_{X}$ respect the Lie algebra structure of real vector fields and that $\sum_{a}P_{a}^{2}$ reproduce, up to a possible scalar curvature term, the Laplace–Beltrami operator on $(\Sigma,h)$. With this choice, the $P_{a}$ generate the infinitesimal unitary action of the isometry group on $L^{2}(B_{\Sigma}(p,r),d\mu_{h})$ and provide the natural, frame-based generalization of the canonical momenta used in flat-space quantum mechanics.
Viewed as operators on the Hilbert space $L^{2}(\Omega,d\mu_{h})$, the $P_{a}$ in (3) are symmetric on the natural Dirichlet form domain (for instance $H^{1}_{0}(\Omega)$ on a domain $\Omega\subset\Sigma$): for all $\varphi,\psi$ in that domain one has $\langle\varphi,P_{a}\psi\rangle_{L^{2}}{=}\langle P_{a}\varphi,\psi\rangle_{L^{2}}$. For a normalized state $\psi$ with $\|\psi\|_{L^{2}}{=}1$ we write $\langle A\rangle{:=}\langle\psi,A\psi\rangle_{L^{2}}$ and collect the momentum components into the vector $P{:=}(P_{a})_{a=1}^{n}$ with mean $\langle P\rangle{:=}(\langle P_{1}\rangle,\dots,\langle P_{n}\rangle)$. In the present geometric setting we measure the distance using the pointwise $h$-norm, and for the canonical momenta $P_{a}$ we introduce the naive momentum variance of $\psi$ by
$\mathrm{Var}^{\mathrm{naive}}_{\psi}(P):=\big\langle(P-\langle P\rangle)^{2}\big\rangle\equiv\sum_{a=1}^{n}\left(\langle P_{a}^{2}\rangle-\langle P_{a}\rangle^{2}\right).$ (4)
Motivated by the structure of the quadratic form of the momenta on $H^{1}_{0}(B_{\Sigma}(p,r))$ (see the integration-by-parts computation in Appendix A) we have
$\langle P^{2}\rangle=\hbar^{2}\|\nabla^{h}\psi\|_{L^{2}}^{2}-\hbar^{2}\,\langle V\rangle$ (5)
with
$V(x):=\frac{1}{2}\sum_{a=1}^{n}X_{a}\big(f_{a}\big)(x)+\frac{1}{4}\sum_{a=1}^{n}f_{a}(x)^{2},\qquad x\in B_{\Sigma}(p,r).$ (6)
and $f_{a}:=\mathrm{div}_{h}X_{a}$ for the divergence of the frame fields. By construction, this potential captures exactly the curvature- and frame-dependent deviation of the geometric kinetic energy from the scalar Dirichlet energy. It is convenient to isolate the purely geometric contribution to the naive variance (4) in the form of a scalar potential [26, 27].
Definition 2.3.
(Geometric momentum variance) Given a normalized state $\psi\in H^{1}_{0}(B_{\Sigma}(p,r))$, we define its geometric momentum variance by
$\mathrm{Var}_{\psi}(P):=\big\langle(P-\langle P\rangle)^{2}\big\rangle+\hbar^{2}\langle V\rangle,$ (7)
that is, by adding to the naive variance (4) the geometric correction term determined by the potential $V$ in (6).
Remark 2.4 (Operational distinction).
It is important to distinguish the physical roles of the two variances defined above. The naive variance $\mathrm{Var}^{\mathrm{naive}}_{\psi}(P)$ in (4) corresponds to the raw statistical spread one would obtain by measuring the frame-dependent operators $P_{a}$. As shown in (5), this quantity mixes the kinematic fluctuations of the state with the local curvature energy encoded in $\langle V\rangle$. By contrast, the geometric variance defined in (7) effectively filters out this frame-dependent potential. It isolates the intrinsic kinetic energy contribution associated with the scalar Laplace–Beltrami operator. Consequently, the lower bound derived in Theorem 3.1 constrains the kinetic capacity required for strict localization within the geometry, invariant under the choice of the orthonormal frame.
In order to relate the geometric momentum variance to the quadratic form of the momenta, we first rewrite the naive variance (4) in terms of the kinetic-energy form of $P$,
$\mathrm{Var}^{\mathrm{naive}}_{\psi}(P)=\langle P^{2}\rangle-\|\langle P\rangle\|_{h}^{2}\,.$ (8)
Inserting the quadratic-form identity (5) into (8) with (7) this yields the intrinsic variance decomposition
$\mathrm{Var}_{\psi}(P)=\hbar^{2}\|\nabla^{h}\psi\|_{L^{2}}^{2}-\|\langle P\rangle\|_{h}^{2}.$ (9)
@@
In what follows we measure momentum uncertainty by the corresponding standard deviation
@@
Remark 2.5.
Physical interpretation and small-ball limit. To clarify the physical significance of the potential $V$, consider its asymptotic behavior on small geodesic balls $B_{\Sigma}(p,r)$. In Riemann normal coordinates centered at $p$, the metric is Euclidean to first order, but the volume element expands as $\mathrm{d}\mu_{h}=[1-\frac{1}{6}R_{ij}(p)x^{i}x^{j}+O(r^{3})]\,\mathrm{d}^{n}x$. The divergence of the frame fields is determined by the logarithmic derivative of the volume density, yielding $f_{a}\approx\partial_{a}\ln\sqrt{|h|}\approx-\frac{1}{3}R_{ab}(p)x^{b}$. Substituting this into definition (6), the quadratic term $f_{a}^{2}$ vanishes to leading order, while the derivative term contributes $\frac{1}{2}\sum_{a}\partial_{a}f_{a}\approx-\frac{1}{6}\sum_{a}R_{aa}(p)$. Thus, for states highly localized near $p$, the expectation value behaves asymptotically as
$\langle V\rangle\;\sim\;-\frac{1}{6}\,\mathrm{Scal}(p).$ (11)
This identifies the geometric potential $V$ as the intrinsic curvature correction – specifically the scalar curvature term $\frac{1}{6}R$ – that naturally arises in the quantization of systems on curved manifolds to maintain general covariance [26]. Operationally, $\langle V\rangle$ measures the energy shift between the square of the Hermitian momentum operators and the covariant Laplace–Beltrami operator.
@@ 3 An intrinsic Heisenberg-type lower bound
From a physical standpoint, Dirichlet data $\psi|_{\partial B_{\Sigma}(p,r)}=0$ idealize a state $\psi$ that is strictly confined to the ball-region $B_{\Sigma}(p,r)$, for instance in an infinite potential well or as the post-measurement state after a sharp localization to that ball: the probability of finding the particle on the boundary or outside $B_{\Sigma}(p,r)$ is zero. Neumann conditions $\partial_{n}\psi=0$ instead describe perfectly reflecting boundaries with vanishing normal probability flux, while Robin conditions $(\partial_{n}\psi+\kappa\psi)=0$ can model partially reflecting surfaces or boundary interactions localized at $\partial B_{\Sigma}(p,r)$. The intrinsic lower bound proved below thus pertains specifically to this idealized regime of strict confinement modeled by Dirichlet data; more general boundary conditions would require incorporating the corresponding surface terms into the uncertainty estimate and lead to a different physical problem.
Accordingly, in the following we impose Dirichlet boundary conditions throughout and work with the Dirichlet Laplacian on bounded domains. By the Dirichlet Rayleigh–Ritz principle (see [14], Ch. 3), the first Dirichlet eigenvalue $\lambda_{1}(\Omega;h)$ of $\Delta_{h}$ on a bounded domain $\Omega\subset\Sigma$ is characterized by
@@
$\sigma_{p}(\psi)\;\geq\;\hbar\,\sqrt{\lambda_{1}\!\left(B_{\Sigma};h\right)}.$ (13)
Equality holds if and only if $|\psi|$ is a first Dirichlet eigenfunction of $\Delta_{h}$ on $B_{\Sigma}(p,r)$ and the phase gradient $\nabla_{h}(\arg\psi)$ is constant almost everywhere on $B_{\Sigma}(p,r)$ with respect to the probability measure $|\psi|^{2}\,d\mu_{h}$.
@@
Since $\psi=u\,e^{i\phi}$ with $u=|\psi|\geq 0$ and real-valued phase $\phi$, differentiate in an $h$-orthonormal frame $\{X_{a}\}_{a=1}^{n}$ to obtain
$X_{a}\psi=e^{i\phi}(X_{a}u+iu\,X_{a}\phi),\qquad a=1,\dots,n.$
Hence
$|X_{a}\psi|^{2}=|X_{a}u|^{2}+u^{2}|X_{a}\phi|^{2}$
and, using the pointwise norm (1),
$\|\nabla_{h}\psi\|_{h}^{2}=\sum_{a=1}^{n}|X_{a}\psi|^{2}=\sum_{a=1}^{n}\bigl(|X_{a}u|^{2}+u^{2}|X_{a}\phi|^{2}\bigr)=\|\nabla_{h}u\|_{h}^{2}+u^{2}\|\nabla_{h}\phi\|_{h}^{2}.$
This pointwise decomposition is the (local) Madelung identity. It holds first for smooth $\psi$, and by density extends to $\psi\in H^{1}_{0}(B_{\Sigma}(p,r))$. Inserting it into the variance formula (9) yields the exact decomposition
@@
Since $\|\psi\|_{L^{2}(B_{\Sigma}(p,r),d\mu_{h})}{=}1$ and $u{=}|\psi|$, the weight $u^{2}\,d\mu_{h}{=}|\psi|^{2}\,d\mu_{h}$ corresponds to the quantum mechanical Born probability measure on $B_{\Sigma}(p,r)$ introduced at the end of Sec. 2.1. Applying the Cauchy–Schwarz inequality with respect to this probability measure, we obtain
$\sum_{a=1}^{n}\Bigl(\int_{B_{\Sigma}(p,r)}u^{2}X_{a}\phi\,d\mu_{h}\Bigr)^{2}\;\leq\;\int_{B_{\Sigma}(p,r)}u^{2}\|\nabla_{h}\phi\|_{h}^{2}\,d\mu_{h}.$ (16)
Thus the last term in (15) is bounded above by $\hbar^{2}\int u^{2}\|\nabla_{h}\phi\|_{h}^{2}\,d\mu_{h}$, and inserting this estimate into (15) yields the lower bound
$\mathrm{Var}_{\psi}(P)\;\geq\;\hbar^{2}\int_{B_{\Sigma}(p,r)}\|\nabla_{h}u\|_{h}^{2}\,d\mu_{h}$ (17)
@@
Here and in the sequel, “a.e.” (“almost everywhere”) is always understood with respect to the relevant underlying measure; in particular, $\nu$-a.e.” refers to the probability measure $d\nu=u^{2}\,d\mu_{h}$ on $B_{\Sigma}(p,r)$. Invoking (1), the difference of expectations simplifies to
@@
In (21) the fluctuation term is written as a sum of variances of the scalar fields $X_{a}\phi$ with respect to the probability measure $d\nu$. One may rewrite this sum as the squared $h$-norm of a centered gradient vector and then take the $\nu$-expectation. Concretely,
@@
for $\psi=u\,e^{i\phi}$. Hence, equality in (18) holds precisely when the gradient $\nabla^{h}\phi$ is $\nu$-a.e. constant. This is the phase part of the equality condition in Theorem 3.1; together with $|\psi|$ being a first Dirichlet eigenfunction it yields equality in (13). By the Dirichlet Rayleigh–Ritz characterization (12) (see [14, Ch. 3]), for any $0\neq u\in H^{1}_{0}(B_{\Sigma}(p,r))$,
@@
Taking $u=|\psi|$ with $\|\psi\|_{L^{2}}=1$ (hence $\|u\|_{L^{2}}=1$) and combining (18) with (23), we obtain the chain
@@
Taking square roots gives the desired estimate $\sigma_{p}(\psi)\geq\hbar\sqrt{\lambda_{1}}$, i.e. (13). $\square$
Remark 3.2.
Coordinate invariance and kinetic-energy window. The estimate is purely intrinsic to the Riemannian data $(B_{\Sigma}(p,r),h)$: neither coordinates nor the ambient foliation play any role in its formulation or proof. The kinetic-energy window follows directly from (9) and the exact decomposition (22). On the one hand, (9) reads
$\mathrm{Var}_{\psi}(P)=\hbar^{2}\|\nabla_{h}\psi\|_{L^{2}}^{2}-\|\langle P\rangle\|_{h}^{2}\leq\hbar^{2}\|\nabla_{h}\psi\|_{L^{2}}^{2},$
since $\|\langle P\rangle\|_{h}^{2}\geq 0$. On the other hand, the expectation in term in (22) is nonnegative, so
$\mathrm{Var}_{\psi}(P)\geq\hbar^{2}\|\nabla_{h}u\|_{L^{2}}^{2}=\hbar^{2}\|\nabla_{h}|\psi|\|_{L^{2}}^{2}.$
Taking square roots in the two-sided inequality for $\mathrm{Var}_{\psi}(P)$ gives
@@
which is the claimed kinetic-energy window. Equality at the upper endpoint holds precisely when the mean momentum vanishes, $\|\langle P\rangle\|_{h}=0$ (equivalently, $\langle P\rangle=0$), so that the subtracted term in (9) disappears. Equality at the lower endpoint holds precisely when the fluctuation term in (22) vanishes, i.e. when $\nabla_{h}\phi$ is constant ($\nu$-a.e.). If this lower-endpoint equality holds but $|\psi|$ is not a ground-state Dirichlet eigenfunction, then the Rayleigh quotient is strict and the bound (13) is correspondingly strict: $\sigma_{p}(\psi)>\hbar\sqrt{\lambda_{1}}$.
Remark 3.3.
On the ‘constant phase gradient’ condition and frame independence. By the exact variance decomposition (22), equality at the lower endpoint of the kinetic-energy window (25) (equivalently of (18)) holds if and only if the $\nu$-variance of the phase gradient vanishes:
@@
Realizability / integrability. Writing $v:=\sum_{a=1}^{n}c_{a}\,X_{a}$, the condition above requires $\nabla^{h}\phi=v$. Such a phase exists if and only $\nabla_{i}v_{j}-\nabla_{j}v_{i}=0$, for all $i,j$. On curved balls this typically fails; hence equality in (13) is often not attained (the bound is sharp but strict). In the Euclidean ball $B_{r}(0)\subset\mathbb{R}^{n}$, however, $\nabla_{i}v_{j}=\nabla_{j}v_{i}$ for constant $v$, and affine phases $\phi(x)=k\!\cdot\!x$ realize the lower endpoint together with the Dirichlet ground-state modulus.
Remark 3.4.
Example saturating the bound with nonzero mean momentum. Let $B_{r}(0)\subset\mathbb{R}^{n}$ carry the Euclidean metric $h$. Let $u_{1}>0$ denote the $L^{2}$-normalized first Dirichlet eigenfunction of $\Delta_{h}$ on $B_{r}(0)$, with eigenvalue $\lambda_{1}(B_{r}(0);h)=:\lambda_{1}$. For fixed $k\in\mathbb{R}^{n}$ set